The APOcrypha

Volume 80, Issue 6 Wednesday, April 20, 1994 Contents

Editor’s Sigh of Relief 1 Candidate Articles! 1

President: Kate Curtis . . . . . . . . . . . . . . . . . . . . . 1 President: Bridget Spitznagel . . . . . . . . . . . . . . . . . 4 Membership Vice President: Dave Cho . . . . . . . . . . . . 7 Administrative Vice President: Gilbert Leung . . . . . . . . 11 Treasurer: Henning Colsman-Freyberger . . . . . . . . . . . 12 APOcrypha Editor: Bert Dvornik . . . . . . . . . . . . . . . 12 APOcrypha Editor: Ping Huang . . . . . . . . . . . . . . . 13

PTRM! 15 President’s Corner 15 What I did for spring break… 17 Just another article. :) 17 Pledge Article: EGW 18 A Random Article… 19 Henning’s Random Ramblings III 20 Pledge Talk 20 Silly Stuff 22 Patriots Day Weekend: The Editor’s Journal 22 Project Report: 1994 Institute Screw 23

The Preparation . . . . . . . . . . . . . . . . . . . . . . . . 23 The Voting and the Results . . . . . . . . . . . . . . . . . . 24 The Crew . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

CPPC Report 25 Stuff from mrn: Random News on AX Alums 29 In Memoriam: Earle M. Herbert 29 Chapter Meeting Summary: April 6, 1994 30 ExecComm Summary: April 13, 1994 32 More Spring Breaking 33 More Dave 34 And One More Dave 35 Advertisements 36

Editor’s Sigh of Relief

Whew! Mucho thanQs to everyone who helped keep this issue of the APOc from being one sheet of folded paper! In fact, it probably broke several records! I may have to write an article about it! :)

YiAPOc, Amy

Candidate Articles!

President: Kate Curtis

I am running for president, because I think that I can continue to motivate the chapter, while giving ExecComm the direction and help it needs to implement the service/membership/fellowship/whatever programs for the fall. This article is really a collection of random things I’d like to see happen (not in any particular order); please let me know what you think, whether or not I get elected. :)

Summer: I’d like to see a couple of service projects and lots of random fellowship, from dinner trips to ???? And, of course, July 4th on the Esplanade. For those of you who are not going to be here for the summer, I hope that you will miss lots of fun, and we look forward to seeing you again in the fall.

Lots more service: I like going on service projects. (That’s probably why I’m still hanging around APO.) Hopefully you all do too. I’d like to ensure (or help the SVP ensure) that as our chapter grows, the number and type of projects change accordingly. We may have the manpower to run double projects some weekends (Saturday and Sunday), especially if some projects can only take up to ten or so people. I also want to ensure that the service program has a large number of service projects in the first third of the term when people have lots of free time. This will also give pledges lots of opportunities to meet brothers. To do this, the SVP is going to need an active ServComm. We need a ServComm to nurture future SVPs too. Really, it’s not that hard a job, and it’s fun. :)

Lots of pledges? Well, I hope so. Most importantly I’d like to see them becoming a part of the chapter more quickly. They don’t have office keys so they can’t hang out in the office, unless, of course, lots of brothers are already there. :) At CPPC we discussed communication channels between the chapter and the pledges, or, rather, lack thereof. I hope to help the MVP make the improvements suggested at CPPC a reality. The idea is to make sure that the pledges feel a part of the APO community as quickly as possible and to make them feel more “clued-in.”

Chapter Skills: Holding another chapter skills weekend is an option. I’d certainly like to continue the push to revive silkscreening and train press ops to the point where the skills have enough momentum to keep themselves going. I would like to see both the silkscreen room and the press shop used extensively over the summer.

$ Money $: The first thing I’d like to do is have someone (maybe me, maybe not) go talk to the registrar’s office. Maybe we’ll get some useful information like what they see happening to Reg Day in the next few years. Then we need to analyse what we are spending money on (sounds like a job for the new treasurer) and figure out where we could best curtail our spending. We spend about $4000 a year; do you know what we spend it on?

Inter-Chapter Stuff: Yeahh for Sectional Service Project. Maybe we could organise a service project between all the chapters in Boston; it shouldn’t be too hard. Conclave can be fun; come and keep me company there.

Academics: Always come first. If anyone has an idea of a thesis topic for me, preferably one that has something to do with databases and multi-media educational initiatives/tools, please let me know.

Fellowship: PARTY!! Well, not just that, that and much more. I like study breaks. I like some of the more innovative breaks that I’ve heard of in the past, like study breaks in elevators with popcorn. Spontaneous fellowship is always good; be spontaneous. The fellowship program is also a way of making the pledges feel more at home.

Nationals: Dallas in December. Try forgetting about how much it costs to fly to Dallas and thinking instead about whether you would like to represent the chapter at the national convention. Do you want see first hand what happens to the motion (about blindfolding during ceremonies) that we are proposing? The chapter maintains a National Layaway fund that can be used to defray the expenses of going to Nationals, so don’t just dismiss the idea out of hand. (On a personal note, seeing as how Nationals is between Christmas and New Year, I will be in Dublin and not in Dallas).

Meetings: Should always be short, interesting and well-run. I can try. :) People seem to like knowing that meetings will not normally run over an hour. I like meetings where something worthwhile is accomplished; otherwise, spending even an hour on it seems too much. I would also like to differentiate better between chapter meetings and ExecComm meetings, although right now I’m not sure what belongs where… Brothers should be equally welcome at either, but they are intended to serve different needs.

ExecComm: Should work together to make the term a success for all the brothers and pledges and alumni and other people. Regardless of responsibilities of the individual office, each person’s general contribution to ExecComm is very important. I’d also like to see lots of new faces, together with the old (not just agewise) ones.

By-laws: As a personal goal I would like to get together a small group of people interested in amending the by-laws so that they are consistent and practical. For example, the references to the budget are inconsistent and the by-laws have no escape clauses for when you try to elect someone for an office for which there are no candidates. The recent national by-laws discussions also raised the question about inserting an anti-discriminatory statement in our by-laws, since it was removed from the national by-laws. I am sure that many more issues will come to light when we start discussing them.

Alumni: They’re nice people, but they’re sometimes a little confused about what the chapter needs from them, or wants from them, or thinks it needs or wants, or just thinks in general. Do we want a Big Uncle sort of program? Do we care? Why did the idea not get implemented anyhow? Were we lame? Were they lame? Were we all lame? (Probably) I can try and talk to the alumni association to see what’s happening.

Finally…

In answer to Ezra’s favourite question: I haven’t a clue what I will be taking next term. I’m a grad student. I have an RA. I will be living in e40 so I don’t get to the office more than once a day (this is a lot less than it used to be. :) Some days I don’t visit the office at all :)). I will be probably be taking Japanese III and a random management class while working principally on my research.

Reason for running now: I thought this term was great; I like the higher level of enthusiasm and fellowship that Dave has inspired in the chapter. Dave has/had plans for the chapter, some of which still need to be nursemaided a little. I can do this while implementing some of my own ideas or other people’s ideas. I get the whole summer to motivate myself and everyone else and to get ExecComm together and get organised. The chapter succeeded in its goal of getting a large pledge class, but now we are faced with the challenge of turning this into an increase in the number of active brothers.

Reason for not running now: A term off ExecComm would probably be good. For those of us who are on ExecComm far too often, we often forget that we see the chapter very differently from those brothers who are less involved in decision-making and planning. It always feels good to be “just a brother.”

Kate

President: Bridget Spitznagel

This is about the first time I’ve remembered to write a candidate article before the relevant elections meeting, so I’m really going to give the APOc its money’s worth. :-)

Hi. I’m Bridget and I’ll be running for president this evening :) For your amusement I may be running for some other offices afterward. I would wear my Australian hat again, but I lost it earlier this month (wah).

What do I think the president does?

The official duties of the prez involve various kinds of coordination. (Fortunately, no hand-eye coordination.) To make a silly paraphrase, each officer and committee within the chapter deals with a different subset of the things necessary to keep the chapter running in the present and to prepare it for the future; it’s the president’s job to make sure the separate parts are working together and to keep the whole collection from running off the side of the road into a tree. The president also makes sure the chapter stays in contact with advisors and with the rest of the world. Lately the prez also gets the job of talking about leadership at open meetings and thinking about holding leadership workshops (if I get organized, I might do more than think about it ;-) ).

What issues do I think the next president should be concerned with?

There are several changes likely to occur in the next few years (some more abruptly than others). The next president should, as always, be prepared for probable changes and be ready to deal with the problems they are likely to cause. She/he should also consider aspects of the chapter that could be changed for the better, suggest these possibilities to the relevant officers or committees, and coordinate and assist their efforts if necessary.

You may have noticed that some people in the chapter have gotten a bit nervous about the future, particularly in reference to the budget. Strictly speaking, the budget is the treasurer’s job and is usually prepared by examining the previous few budgets and deciding whether they were reasonable or not. (Note that “reasonable” and “realistic” are not necessarily related concepts. The realistic budget is a rare beast.) The general consensus among the nervous people is that the ’Tute is going to computerize Registration Day completely, thus eliminating our major source of income; however, there is no real agreement about exactly what we should do about this when it happens, nor about whether we should start doing something about it before it happens. In other treasury matters, security has recently become an increased concern.

You may also have noticed that the chapter is getting larger and that there are more people running for offices (yay :-)). These are Good Things (tm), but a good thing can still cause problems if it is not planned for. One of the immediately apparent difficulties that an increase in membership may cause is lack of communication. When there are around twenty active brothers in the chapter (roughly half of them are officers and attend meetings regularly), communication is barely an issue; word of mouth is nearly sufficient. A larger chapter, assuming some diversity in majors and living groups (e.g., not everyone lives in New House), is likely to require more organized lines of communication. E-mail makes this less of a problem than it might be, but should not be relied upon entirely. The service program itself would have to grow in some direction (either larger projects, or more of them) to maintain current standards of per-brother service, and the recruitment and membership retention programs would have to grow to maintain the size of the larger chapter. (Actually, there is currently no membership retention program.) As for the minor issue of an increased number of people running for offices: In the (fairly recent) past, elections consisted of finding a candidate for each office and electing her/him. Now that we have a greater opportunity for choice, we will have to decide what criteria are actually appropriate (for example, should one person be favored because this is the last term she/he can reasonably be expected to hold this office? should… never mind, you can come up with weird examples on your own). An increased size may affect leadership in other ways as well (e.g., people may become more project-specialized, or officers might run projects only rarely.)

Our relations with other groups are likely to change as well. We are beginning good relations with the chapter at BU; there is another service group (Circle K, largely Baker-based) on campus which has expressed interest in establishing occasional relations; people are trying to create the tradition of a Sectional Project; more people in the chapter have started attending conclaves (yay). The chapter’s selfimage (how members perceive AX itself) is likely to affect the way that we see and interact with (or don’t interact with) outside groups.

So what am I doing next term, anyway?

(This new VI-3 degree program is confusing :-S. I used to have everything planned out… then they changed everything. Cool! techinfo has the course catalog online! … but not degree requirements.) My academic commitments will be fairly serious (though nothing to really worry about). Call it Japanese III for a HASS-D, and two or three CS classes (depending on which ones they are.)

Non-academic commitments: If I am elected to a major office, I will not participate in any organized musical endeavors. (They seem to be specially planned to conflict with APO.) It is unlikely that I will be able to find a paying UROP. If given the opportunity (the fall class is smaller, so they don’t hire as many) I will lab-assist 6.001 again; that takes about eight hours a week arranged any way I want. Thus, I should be able to attend all meetings (and open meetings, and pledge ceremonies, provided they are planned at least a week in advance.)

What else?

If not elected president, I may run for APOcEd (I forget who’s running already.) I know LaTEX, and I would put out APOcryphae regularly (and probably even remember who to mail them to :) ). I can’t guarantee content (you can only twist arms so far before they break off and their owners can’t type anymore), but if I didn’t get any actual content I might stick in some useless but amusing things, such as a pack of lies column, or a how to cook unusual things in a microwave column; or I might reprint interesting/useful bits of the pledge manual, the chapter book, how-to documents, etc.

If inadvertently elected Fellowship Chair, I promise to be interesting (you can take that as a promise or a threat; either way).

YiLFS, bridget

Membership Vice President: Dave Cho

I’m running for MVP. This article addresses the following topics in the following order:

1. Who am I? A brief introduction.

2. Why

• Why do I want to be MVP?

• Why I shouldn’t be MVP.

3. What would I do?

4. How would I accomplish my goals?

5. When will I have accomplished these goals by?

6. Where would I hold events?

Who. My name, just in case any of you forgot, is dave cho. I’m currently a first-term senior in VI-3. I first pledged in Spring 1989 and was elevated in Fall 1989. Since that time, I’ve been variously involved with the chapter, most notably as a uniquely lame Fellowship Chair in Spring 1993 and now currently as President. I plan on graduating this December.

Next term, I’ll be taking 6.044 (? Al Drake’s algorithms course, which I hear isn’t terribly hard), two music courses (yay!), and thesis.

Why. Why do I want to be MVP? Well, a little history. Way back when I was a pledge (both terms), I said to myself, “Someday I want to be MVP and President.” No, not simultaneously. So this isn’t a whimsical thing. Nor is it because the slot is empty. Nor is it because I think the other candidates are lame (of course, I know of no other official candidates). It’s because I really want to be MVP. I’ve seen a lot of MVPs, and their different styles, and the problems they’ve addressed. I’ve worked on the pledge program by helping pledges formulate their problems; I’ve worked with specific brothers in identifying and addressing problems; I’ve worked with MVPs directly. But I haven’t done the job personally, and I would like to.

I’m old. I’ve seen things the way they’ve been done before. I understand the chapter, or at least, I think I do. :) I have confidence in myself, in my abilities, in my understanding of things.

I love this chapter, and the people in it, and what we do. That’s why I ran for President: because I was interested in doing more, in making the chapter grow, in making this chapter better. And a little secret plan I’ve had for awhile was to be Prez and “set things up” for a really keen term as MVP. So I’ve been trying to get things into order; to get brothers used to the idea of greatness. And the logical followup is to bring a pledge class to take up the work where I left off (oh God, this is really stupid, but i’m tired and it’s 12:37 am, so please bear with my inane writing). But it’s a dream, and I’m working on it, so there. :)

Now, before anyone gets a really stupid idea, of COURSE I’ll understand if you flush me like a dead fish. Don’t be silly. :) Besides, it just means I’ll have time to do Concert Choir next term, and I hear we’re doing something really cool, and you should come to our concert THIS FRIDAY AT 8 PM IN KRESGE; TIX ARE ONLY $2 IN ADVANCE SO LOOK FOR OUR BOOTH IN LOBBY 10.

Oops, sorry. That shameless plug just happened to slip in.

Why shouldn’t I be MVP? Well, I’m old. There’s something to be said for MVPs being around after their class is elevated. However, I’m not planning on disappearing from Boston. (Yes yes yes, I realize the best-laid plans never materialize, but I’ve avoided leaving Boston for six years, and things as they are certainly point to my being around Boston.) And it is true that Being Outside will preclude a great deal of personal involvement. But I hope to be at least marginally involved, and after the first two months, I don’t expect I will have a realistically significant effect.

(No, really, I’m not tired. Yes, this is a non-sequitur.)

Other reasons not to elect me. Umm… oh yes. I might be so overly sarcastic that I’ll have two pledges elevated, four continued, and eight in a mental trauma hospital. Don’t worry about it. Or that I’ll flirt with all the pledges. Umm… hmm, well, I won’t say that I’ll never flirt with them (I’m being half-facetious, half-serious here) until I at least know them (c.f., this term’s officers and pledges).

In other words, I can’t think of any other reasons not to elect me other than other superior candidates.

What would I do? I think CPPC brought up a lot of good ideas. Umm, and not wanting to be my usual superior self, some of them were ideas I had heard of and/or came up with a long time ago. (Not at all surprisingly, I had since forgotten them.)

Down to specifics:

• BBs by Pledge Ceremonies. This is tricky.

• More formalized Pledge Meetings. I think I would try to have them just before or just after ExecComm meetings. What actually goes into them would be a summer project, but it would certainly be ruled by “what do the pledges want to know and need to know at this point” more than anything else.

• I probably wouldn’t have any other pledge-specific meetings, since realistically speaking, neither they nor I have the time for them.

• Pledge Manuals at Pledge Ceremony Receptions (see CPPC report).

• Pledge Study Break, preferably early in the term. Get them to organize a study break. (Ooooh, ahhh, really hard there.) The point is not a leadership exercise (except for the PC), but rather, an opportunity for pledges to meet brothers and vice versa. In addition, to also get used to working together: a not-so-subtle setup for Pledge Project.

• I think it’s policy, but if it isn’t, I want the Pledge and Voting policy to be voted on at I-Banquet.

How would I accomplish my goals? The Big Brother Network doesn’t work; it’s too big. Doing it along would be a lose. Thus, I’d be getting a MemComm. It’d be á la ServComm. Three or four people. I’d be dispatching events as they came in and making final decisions, but otherwise, working closely with these people in keeping in touch with the pledges. Now, obviously, if all the Big Brothers are angelic, getting eight hours of sleep a night, their weekly netrek game, and three balanced meals daily, things will obviously be OK. But if they don’t, well, that’s why I want to be MVP, so i can play their netrek games for them.

Where would I hold events? What a ludicrously stupid question. It was obviously constructed to have a “where” question.

Closing comments:

1. Yes, I know about the time commitments.

2. Yes, I can be serious. But I think humor is a good thing for an MVP, and while not everyone can appreciate a netrek joke, it still is a dumb joke, and as we all know, dumb jokes are always the not-quite-right form of humor for any situation.

3. I really want this job, so don’t ask me that.

4. I know I’ve missed a lot of important information, but since I won’t be at a 6.004 test tomorrow night, you can just ask me.

dave dave cho MVP candidate

Administrative Vice President: Gilbert Leung

I somehow have succumbed to the charm of Amy and started writing this article that you are reading. What is this article about again? Hmm… right! I’m running for AVP! I’m psyched! I’ll have time! Vote for me!

What? You are not convinced? Ok. I’ll tell you a reason or two why I want to run.

1. I want my name added to apo-acl, the real power! (In case you don’t know, apo-acl is the athena list that controls all the apo mailing lists as well as the apo locker. Once you are on the list, you will have the permission to nuke everything in the locker. >:) ok. ok. That’s not exactly what I’m going to do…)

What? That is still not convincing enough? Hey, when I said a reason or two I meant it! So did I when I said that I’m psyched. Hmm… Voters are certainly more and more demanding everyday… ok. ok. Let me make up, err… tell you another reason:

2. I never really pay 100% attention at meetings (which I consider bad). If I become the AVP, I will have to take minutes and will force myself to pay heed (which I consider good). Or, even better, I shall learn how to pay only 50% attention and still manage to take minutes. :)

So, I have cited my two reasons and don’t you make me contradict myself by asking me to give you another!

Gilbert

Treasurer: Henning Colsman-Freyberger

I am running for Treasurer.

1)

“class”  units
11.001J Intro to Urban Design and Development  3-0-9
11.011 The Art and Science of Negotiation  3-0-9
14.24 Economics and Law  3-0-9
14.54 International Trade  4-0-8
Fencing Team  8-2-0
German House Steward  0-2-2
New House ComCom Chairperson (pending election) 0-3-1
New House Vice President (pending election)  .5-2-1
SafeWalk  0-2-0

Total: 21.5-11-39 = 71.5 units

2) a) I have undergone Treasury training and feel I can competently carry out the duties of this office.

2) b) I have undergone Treasury training and feel I can competently carry out the duties of this office.

Henning Colsman-Freyberger

APOcrypha Editor: Bert Dvornik

I haven’t quite decided whether to run for APOcEd or not. However, I’m going to write and submit an article. I can always decide to back out before [or during] elections.

I have never done any newsletter editing before, but I like to play with LaTEX, and I believe I can learn the tricks of the trade that I don’t know already. Anyway, editing the APOcrypha seems like a fun job.

I’m going to make only one campaign promise: I’m going to be getting APOcs out a day or two before the chapter meeting. Thus, the brothers will be able to read anything relevant to the meeting before the meeting itself.

For Ezra’s benefit, I’ll admit that I don’t really know what classes I’ll be taking the next term. I will have a lot of humanities, though, because I kind of need to catch up on my HASS requirements. I might or might not also be taking 8.07. I think I’ll decide after Junior Lab ends… Without the shadow of Junior Lab looming over my head, I’ll also be able to get more active in the chapter in general.

I can’t promise that I’ll do a better job than any of the other candidates would. If I get elected, I’ll do my best. Since you are the ones who’ll have to read the APOc, I urge you to vote for whomever you think will do the best job, even if it isn’t me. =)

YiLFS and 8.14, bert =)

APOcrypha Editor: Ping Huang

Hello. I am running for the office of APOcrypha editor. Per request (or more accurately, begging^{1}) of the current APOcrypha editor, I am writing a candidate’s article. :)

I will start off by answering the standard questions that Ezra seems compelled to ask every single candidate for any chapter office, and I will be nonplussed if I get asked these questions during the elections meeting (hint).

• The previous offices I have held in the chapter are (in chronological order): Historian, PubDir, AVP, MVP, and SVP.

• I have not finalized what I will be registered for in the fall, but it will very likely consist of two graduate-level computer science classes, a women’s studies humanities class, and something random. Next term will be my ninth term as a student, and I expect to graduate in May 1995.

• My time commitments might include working part-time off campus or doing some research in an LCS or AI Lab group, in preparation for writing a thesis sometime next year. I am currently involved in the Baker Foundation and the Associate Advising Steering Committee, but I am moving to disengage myself from those organizations now that I am a graduate student. There is probably about an even chance that I will be signed up as an associate advisor.

1I have more to say about this later.

I believe I could do a good job as APOcrypha editor, and it is a job which I believe I would enjoy and have considered running for several times before. I know I can handle the mechanics part of being APOcrypha editor: I am certainly familiar enough with Athena and LaTEX, and as a past AVP I know how to deal with labels, mailings, etc.

However, it takes more than familiarity with the mechanics to make a good APOcrypha editor. I have been very pleased to see the APOcrypha come out on a regular basis this term, and my intent if elected will be continue to put out the newsletter on a regular schedule. At CPPC it was mentioned that the APOcrypha could serve as a more effective communications tool for the chapter if it were to be published several days before each chapter meeting rather than just before, so that it may serve as a reminder to brothers that there is a chapter meeting coming up, and to provide them with a copy of the agenda. I will attempt to adhere to such a schedule. I’d also like to try to work with other officers in figuring out how the chapter can be best served by the various communications media available.

In addition, I would like to see more articles published in the APOcrypha which are not minutes summaries or project reports. To that effect, I plan to target (i.e., harass) specific officers and brothers for topical articles as well as essays on issues which arise about or in Alpha Chi. Although I will not resort to threatening to publish extremist political tracts as filler (as some of you may recall past APOcrypha editors have done) in order to coerce articles from the general brotherhood, I will solicit as best I can in general as well. It is my hope to help change how people feel about the APOcrypha, such that APOcEds will not have to plead for articles in a chapter where numerous brothers will no longer feel the need to complain about lack of effective communication.

I’d also like to see more information in the APOcrypha which is less parochial in nature, i.e., more about what other service-minded students are doing on MIT campus and at other campuses/chapters. I am open to new ideas, and may try to implement some which I’ve been kicking around for a while (such as a quotes section, or a 25 Years Ago in the APOcrypha section, an idea with which readers of the Scientific American magazine will be familiar).

YiLFS, Ping Huang

PTRM!

This is a reminder that Pledge Term Review Meeting (PTRM) is

THIS SATURDAY, APRIL 23

The pledges (once again) are:

Shiu Au  Richard Barbalace Angela Chang
Helen Chen Cathy Coury Andrew Garcia
Hannah Jang Aphrodite Kakouros Juliana Leung
Cris Lin Jeffrey Poore Brad Ricketson
Mike Saginaw Andy Su  Grant Su
Junko Tamiya Emily Wallis  Bryan Watabe

President’s Corner

Hiya. I was lame last time and didn’t write for the APOc. Sorry, Amy. Nor have I written a “what I did during Spring Break.” Soon. :)

A few noteworthy events happened over the last four weeks: 1, Big Screw, 2, the office got broken into, and 3, CPPC.

Big Screw

Big Screw happened. Good work, Henning, and everyone involved. I think perhaps next year, we really do want to have a contest ¿ five days. It’s just too hard for the professors to campaign otherwise; the profs really need to mention it in lecture.

The office got broken into

The following report was written by Kevin; I’ve added a few comments.

You have probably heard some references to the incident involving missing money from the Big Screw contest. This is to update you on everything we currently know about the situation.

Thursday night, after the Big Screw money for that day was counted, Henning (Big Screw PC) placed it in the office strong box, locked it, and returned the key to its place in the treasurer’s desk. Friday around lunchtime, the treasurer’s desk drawer was seen wide open. Friday evening when Henning went to retrieve the money from the strong box, all the bills were missing. It was subsequently determined that a few dollars were also missing from the treasurer’s desk.

Thus, the most likely scenario is that late Thursday night, someone gained access to the APO office, took the money and key from the treasurer’s desk, opened the strong box and took the money, locked the strong box, returned the key to the treasurer’s drawer, and left.

Any additional information you may have should be brought to my attention immediately.

WE INTEND TO TAKE THE FOLLOWING ACTION:

1. The chapter will reimburse the Big Screw contest out of the general service fund. (Recommended at last Wednesday’s ExecComm.)

2. Effective immediately, cash will NOT be left in the treasurer’s desk drawer. It should be placed in an envelope with any accompanying receipts or messages and slid into the strong box, which will be kept locked at all times.

3. The key to the strong box will no longer be kept in the treasurer’s desk, the president’s desk or any other accessible location in the office.

In addition, I remind you that the office door must be kept locked at all times unless someone is present in the office, and that someone must be AWAKE and in the MAIN office, NOT in the back room where the press and the tools are.

Thank you for your cooperation.

YiLFS,

Kevin Dahm APO Treasurer

In addition, at ExecComm, it was recommended that I look into getting the lock replaced (but with the same pin heights) and perhaps a spring for the door (to ensure that the door closes the last few inches).

If you have any further questions, feel free to send me mail.

CPPC

CPPC happened at Chapter Weekend, and was very successful. My report is included outside this writing.

dave President

What I did for spring break…

[Editor’s Note: This article was actually submitted before Issue 5’s deadline, but somehow it got overlooked. My apologies to Ping.]

The weather wasn’t very spring-like. And it wasn’t much of a break. But I did sleep a lot, so that was good.

Ping

Just another article. :)

It’s been a long time since I’ve been involved in an APO project/social event, and, quite frankly, it feels rather funny. (It’s Tuesday morning, just minutes before 3:00 am, so, please, bear with me. :) )

I remember being “fairly” involved in APO in my freshman year. It was really cool to go to projects and paint and/or destroy various inanimate and/or animate objects; it was really cool (though really masochistic) to manage to put together a little Semi-Formal Formal almost alone during one of Boston’s worst blizzards. And the elevation ceremony was something that is not soon forgotten…

Hmmm… why was I inactive this term? I had this belief that I could divest myself of my activities; that way, I could spend more time on my work. Now, how many of you think that actually happened? I thought so. :) There’s this little voice inside my head that keeps telling me that I would have done no better if I didn’t become inactive and managed to make the Semi-Formal Formal (sorry Gilbert… but right now my brain is so burnt out that I can’t even remember why I couldn’t make it…). But I guess no one else here has my GPA…:-/

Well, it’s 3:15. I think it’s about time to get back to my 14.02. I’m almost a half-dozen lectures behind (can you say “6.003”?), and I have an exam this Thursday on that material.

Good night, all!

Daniel

(with only \sqrt{0.2} brain cells left for tonight)

Pledge Article: EGW

Hi! It’s Emily, who has now FINALLY gotten around to writing something for the APOcrypha. Yay!!!!!! I’m a freshman. So, yes, I’m on pass/no-record, but I am still hosed. I’m taking 1.00, yes, that’s right, 1.00, the class whose professor just won the 1994 Big Screw Contest! So, if you will excuse me, I have to go back to laboring on for the ’Tute’s biggest slavedriver this year. No one can ever be done with the work for that class; Logcher was the rightful winner. AAARRRRRRGGGHHH!!!!!

YiLFS, Emily

A Random Article…

Hiya…

I’m just a brother who’s gone incommunicado for a while, all because of classes. :-( I even had to put off volunteering at the hospitals and forego fun chapter projects to keep from flunking a few classes. Yup, I do know a few of the pledges on the list… but I haven’t had a chance to talk to them. Even Mike, whom I’d met during Rush last year. I do hang around the APO office when I get a chance (usually at rather odd hours) … and just recently, I saw how Ping printed some cards on that odd printing press. I’ve resolved to learn printing the very next chance that I get. I even read a children’s book on Hiroshima that was given to APO (I recommend it to everyone… it’s on the APO bookshelf near the historian’s desk). Has anyone compiled the top ten reasons to go to MIT yet? Or better, the top ten reasons for being in APO?

Well anyways, hope to see many of you around (and I’ll try my best to “reactivate” myself). Until then, life is just one indefinitely long night of tooling.

YiLFS, Cindy! :-)

Henning’s Random Ramblings III

The reason I asked last APOc’s trivia question (What is the color of the covers of United States passports currently issued?) is that it served as a reminder to renew mine. This I have done in the meantime, so I am well informed now. The correct answer is dark blue. Congratulations are in order for Ann and mrn, who got it right. I had thought that the color had been changed recently and so I was surprised to be informed that my passport would be dark blue again. After further inquiry, my suspicions were confirmed. The color of the passport had changed about 18 months ago to an olive drab, but had been changed back to the blue a year later.

This week I thought I’d pose three questions:

• One to encourage office exploration: What are the flavors of the sodas sold by the APO Coke machine and what is the price?

• One to promote exploration of Athena: What is the ratio of “No Smoking” to “No Food or Drinks Allowed” signs in the student center cluster?

• One to challenge non-MIT intellect: What is the political status of Greenland?

That should provide enough scope. :) As usual, send your answers to trivia@mit.edu. Also, I am offering a new service. If you have any questions about anything APO-related, feel free to ask them and I will answer them in the next issue of the APOc. Anonymity is guaranteed if requested.

Henning Colsman-Freyberger

Pledge Talk

Well, it is almost the end of the term. A lot of people are getting hosed now, and finals are just around the corner. But as everyone knows, so is PTRM. And soon is the meeting where pledges will be initiated into brothers. I figure that I must restate what I said at CPPC. I know that whatever I said there made people laugh (of course there was a lot of that going on). What I basically talked about was this: I am an Athena nerd (and God help me but I know that I am not alone). This means that I read my mail almost three or four times most days. However, I listened to what was going on in the Chapter BEFORE I pledged. That was how I got a lot of my info. I listened. I didn’t have to ask much. I just knew there was an apo-news mailing list, and I added myself. (I do think I remember asking Barbara and her telling me about the list.) I read the mail, sometimes with no clue what the main point was. But sometimes, I got a lot of info out of the messages. Like what APO is about. What we do. And of course, I pledged, and then everything that came in over apo-news was informative. Now I am an Athena nerd, remember, so this method works for me. I would say for those non-Athena nerds, we need to find a way to convey this info to them. (I am not going to elaborate because this was a big subject in CPPC, and Dave will probably fill a lot of you in.) The other thing I mentioned was some thoughts I have about apo. People might think I am sweet-talking the chapter a little, but I am not being untruthful when I say that APO has grown on me… I have met a whole large number of new friends through APO. I have found a place that I like to hang out. In fact, doing homework at the office is somewhat more convenient especially since my computer is far away… PCing chapter weekend is another report that I will write, and I am sure that you will all see this later. But I must say that I had a good time PCing. Granted, we will do a tool list next time so that we don’t end up with extra tools. But it looked like everyone enjoyed the weekend and I know I did. I must now put in a good word for my fellow pledges. I must say that I still can not believe we have 18 pledges. But a lot of them I do know, and I must say that there are some really great people out there. I would like to thank my PCs for a job well-done: Cathy, Transport PC; Aphrodite, Grant, and Hannah, Food PCs; Junko, who survived her second PCship as my Manpower PC, and Cathy as temporary (Sunday) on-site PC. Thanks to all the pledges and brothers who came to help at CWII. Thanks to everyone for making me feel welcome in APO. I wish everyone luck in the final weeks and of course, good luck to APO office candidates

YiLFS & Pledging, Jeffrey

Silly Stuff

Junko and I were being silly at Chapter weekend…

I always say “and my a– is painted green” when I don’t believe her. Well, we were painting the shutters green, so she decided to paint my butt green too… so that day, my butt was TRULY painted green.

=]

Cris

Patriots Day Weekend: The Editor’s Journal

Well, here I am again, writing another article for myself. I guess I just like to read my words in print. Anyway, this past weekend (it’s now April 19, 7:27 pm was, amazingly enough, productive, relaxing, and sleepful! How is that possible? I have no idea. Time warp, maybe.

Friday: The weather was absolument fantastique! I went jogging. In the evening I watched Groundhog Day, which was pretty funny (one Kleenex, though). Anyone want to venture a guess as to how many times Phil went through that day?

Saturday: The weather was wet. I got soaked on the way to the student center for the Technique meeting. Serves me right for being late. Anyway, for three hours I helped design and paint a poster, getting my socks very dirty and pink-spotted in the process. I also spent another three hours in the darkroom doing living group reprints of Phi Kappa Sigma. Then I went to a lobster bash and stuffed myself. Oof.

Sunday: I slept late! I did three hours of work (envelope-stuffing, counting bubble sheets, tying string around old forms) for CEG, then played three hours of volleyball, then ate dinner at Café of India in Harvard Square. Hey, I don’t know how authentic it was, but it tasted good to me (onion pita bread stuff, meat samosa stuff, lamb and mushrooms stuff drowned in brown sauce, and chicken stuff drowned in orange sauce).

Monday: Slept late again! Read Book III of Paradise Lost and actually understood most of it. Amazing. Napped. Discovered that Domino’s Extravaganza pizza on thin crust doesn’t give me too bad of a grease cough. Napped. Read some 12.001 (a. k. a. “Rocks for Jocks,” but the class is actually a pretty big time sink).

Tuesday: Didn’t sleep late, darn. Worked desk at New House, wrote three letters, delivered mail, gorged myself on Smartfood, sour cream and onion Pringles, and Chewy Chips Ahoy cookies, read more 12.001, napped, worked on the APOc, went to a surprise party for Emily Berkowitz of the Alumni Office (soon to be Emily Sandberg!), had cake, worked on the APOc, went to the Tech news meeting and somehow managed to pick up two stories to write for the next two issues, ate Chinese food, and came back to the APOc… where I still am (April 20, 12:57 am).

YiLFS, Amy

P. S. Final copy of this issue printed a little after 1:55 am. Whew.

Project Report: 1994 Institute Screw

The Preparation

The candidate search began Registration Day and ended the day the voting ended. Numerous potential candidates were approached, both faculty and staff. It is amazing how professionally these people can say no. “Foo is a very sensitive issue right now.” “I’m not that thick-skinned.” But these and other answers did not stop the quest for candidates. Publicity included an LSC slide, a drop poster, flyers, and word-of-mouth advertising. The dexion was set up Sunday, April 3, 1994 (the year for anyone reading this in the distant future :) ).

The Voting and the Results

Votes ($) were cast 10 am-5 pm each day, April 4-8. The final tally:

Robert Logcher 1.00 $179.12
Jacob White  6.003 $138.00
Irwin Oppenheim 5.60  $56.74
Stephen Lippard 5.03  $44.75
John Guttag  6.170 $30.30
Bill Clinton  Hack  $4.96

A total of $492.80 was raised. It will go to the Planned Parenthood League of Massachusetts. New policies concerning the handling of money resulted from the contest [see ExecComm Summary].

The Crew

Candidates PC  Jerry Lin
Manpower PC  Bryan Watabe
Money PC  Shiu Au
PC  Henning Colsman-Freyberger
Publicity PC  Richard Barbalace
Publicity PC  Dima Dribinsky

Many thanks go to the PCs and the many workers who manned the booth. Without all the help, this logistical task wouldn’t have been the success this year’s Institute Screw Contest turned out to be.

YiLFS &

Henning Colsman-Freyberger 1994 Institute Screw Project Chair

CPPC Report

I want to start with something that I feel is really important.

After CPPC, quite a number of people came up to me and said something like “Great meeting” or “That was an excellent CPPC” or compliments along those lines.

Not wanting to sound either unappreciative or overly modest, but I really must respond via:

“Bulls—.”

Let’s get real.

The President does not make the CPPC. Just because I was moderating the meeting does not mean the CPPC was “good” or “better” or anything. Now perhaps I moderated it well. But this wasn’t anyone’s point.

I feel, and I think others would agree, that this CPPC was excellent in identifying problems and starting on solutions. This is good. But this is not because of me.

It’s because of the people who were there, who put in three hours to stay awake, to concentrate on what people were saying, to listen, to think, to try.

CPPC’s success emanated from two things: that a lot of people were there and that those people tried. This is the case with most things in the chapter. Never forget this. Never stop trying. Never stop trying to work with new people. Never stop trying new ideas.

So. Onwards to the rest of my report.

In attendance: Elth Ogston, Hannah Jang, Cris Lin, Junko Tamiya, Charley Hamilton, Ping Huang, Yale Zussman, Kate Curtis, Alan Meisler, Jeffrey Poore, Richard Barbalace, Ezra Erb, bert Dvornik, Cathy Coury, Glenn Berry, Ann Yen, Henning Colsman-Freyberger, Kevin Dahm, dave cho.

Time: 8:45 pm to 12:05 am.

The topics brought up for discussion were • Membership: recruiting

• Membership: retention

• Expansion issues: how the chapter needs to deal with its increasing size

• Relations between officers and the rest of the chapter

• Lack of PCs

Ezra’s summary of the beginning is excellent:

We started with communication between officers and brothers/pledges, with an emphasis on how to make brothers/pledges comfortable with bringing problems and difficulties to the officers. Next it mutated into a discussion of chapter meetings vs execomms, and how to make them distinct. Then it mutated into a discussion of email and email lists. From here it degenerated into trivia, and Dave wisely called a recess.

Kate’s conclusions are insightful:

Discussion about relationship between execcomm and the rest of the chapter was not really very productive except for the bits that suggested that we look at the differences between chapter meetings and execcomms and what people want from them. Short meetings are appreciated, but this should not be at the expense of discussing chapter business or ignoring business which perhaps ought to be raised.

The few concrete ideas from the beginning half of the meeting were:

1. Announce officer hours at Chapter Meetings. Also, publish them on the inside cover of the APOcrypha.

2. At Chapter Meetings, have a one-page handout of upcoming events; basically, publish a new WUAPO for Chapter Meetings.

After some time, the pledges spoke up and mentioned how they weren’t getting the information they needed nor when they needed it. The rest of CPPC was devoted mainly (1:45) to Getting Clue To The Pledges; the last 20 minutes was about Making Pledges Part Of The APO Community. The other topics above were not discussed.

The part about Getting Clue To The Pledges mainly emphasized the point that the brothers are not telling the pledges neither what they need to know, nor when they need to know it. Pledge manuals are far too late; they should be coming out at Pledge Ceremonies, as usually planned. Pledges don’t necessarily know “what’s next,” what they should do next, how they can find out about it, etc. Too much that is automatically assumed by the brothers (the existence of signup sheets, what they’re for, and where they exist) is screwing things up.

Ideas/solutions:

• HTML [Editor’s Note for the Non-Nerdy: HyperText Markup Language] the pledge manual? 1/2 :)

• There are four varieties of contacting people: email, phone, snail mail, and personal. Know the advantages and disadvantages of each.

• The pledge manual needs to have very basic information. Things like:

– signup sheets exist

– email lists exist

– chapter meetings and ExecComms exist; and for this and the above, explaining what they are.

• A mandatory reception after the pledge ceremony is a good idea. During this, feed them ice cream, give them an office tour, give them their pledge manuals, and a pledge manual quickref.

• The quickref should explain very basic things in two ways: what, and how.

• A real big problem for pledges is: What now? What do I do now? Where should I go next? Who should I talk to?

• The tradition of Big Brothers pinning their Little Brothers at Pledge Ceremonies should be reinstated. This will require a pre-Pledge Ceremony meeting, which is a good idea anyways to explain more of the chapter.

• Information density, especially of the office, is a problem. The pledges are being hit with a firehose, and don’t have a context to process it with.

• Everything in the office should be labelled, e.g., “This is the Signup Sheet board. Signup sheets, except when at meetings, always live here, so you can come in any time and sign up for project.”

• Open Meetings need to be restructured again; prospectives still aren’t getting the information they’re interested in. For instance, we should give them a “roadmap” for the future: what they need to do, what meetings they need to come to, how to find out more, etc.

It should be stressed that the current inadequacy with the pledge program is the result of many years of problems. Some of the above problems stem from recent sliding of discipline over the last few terms; some have been prevalent since I was a pledge six years ago.

Onwards.

We finished by discussing the problem of incorporating pledges into the “APO Community.” Suggestions included

• We should explicitly point out that brothers aren’t trying to find reasons to not elevate pledges, but rather, reasons to elevate.

• A concern of many brothers is legacy; in a few years, we’ll be turning over the chapter to these pledges, and we want to make sure that it’ll be in good hands.

• Personally ask for a pledge’s opinion. Telling a group of pledges we want to hear from them does not have the same effect as going up to a pledge and specifically asking them for advice, help, assistance, etc.

That’s it for this term’s CPPC. Many thanks to everyone who came and participated; literally speaking, it couldn’t have happened without you.

Now for the hard part: now for implementing the changes. I wish the chapter and imminently-incoming ExecComm luck with their plans.

dave President

Stuff from mrn: Random News on AX Alums

[Editor’s Note: mrn is our APO Regional Director.]

Recently spotted in Washington, at the rechartering of Central Washington University: Liz Wise and Jock Young, AX circa mid-70s. Jock is an advisor to the new chapter.

And closer to home, Rick Hartley, an AX alum, was elected as Section 99 (New Jersey) Chair this past weekend.

Alpha Chi must be doing something right in the leadership development department. This means every Section in Region I has been chaired or represented by AX alums (and of course, there’s your beloved Regional Director, too. ;) Not to mention Sections in other parts of the country, most notably Region X, and in fact, the participation at all levels of the Fraternity.

mrn

**************

In Memoriam: Earle M. Herbert

It is my sad duty to report that past National President Earle Herbert passed away this past Friday.

To be honest, I am somewhat at a loss for the right words here. Suffice it to say that Earle will be greatly missed, and I am just so happy that my life was touched by this great man. Alpha Phi Omega has lost one of its giants and a man who, more than anyone I know, epitomized our cardinal principles, especially friendship.

Earle’s memorial service will be at 1:30 pm this coming Saturday at St. John Fisher Church, 5448 Crest Road, Rancho Palos Verde, CA 90274.

Expressions of sympathy may be sent to “The Family of Earle Herbert,” 3671 Vinton Avenue, No. 8, Los Angeles, CA 90274. This is Earle’s home address, and his family will be there for several weeks taking care of personal matters.

Both Earle and his family have asked that, in lieu of flowers, donations be sent to the Alpha Phi Omega Endowment Fund, 14901 East 42nd Street, Independence, MO 64055-7347, in the name of Earle and all future pledges. From a financial standpoint, this fund is our future, and Earle always wore his pledge pin because he felt that a true brother never stopped pledging. Thus, such a gift is most fitting and would be the best honor we could give Earle, especially because of his dedication to and respect for pledges.

Personal note: I hate to ask for money. I do it with alumni because that’s part of “the job.” But I have to tell you that even $1 in Earle’s memory to the Endowment Fund will be greatly appreciated and will honor a truly great person who gave everything he had to APO. But I especially ask for your prayers for Earle and his family, and I ask you to be grateful, as I am, for the fact that APO had an Earle Herbert.

Jerry Schroeder National President

                            gschroed@cap.gwu.edu

**************

Chapter Meeting Summary: April 6, 1994

The meeting began at 7:33 pm in the Mezzanine Lounge on the third floor of the student center. People who were there at one time or another: Dima, Ezra, Jo-Anne, LeaAnn, Henning, Dave, Bryan, Kate, Rachel Yale, Richard, Kevin, Charley, Bridget, Gilbert, Cathy, Harry, Ann, Joe, Barbara, Amy, and Ping, not necessarily in that order.

The SVP reported a past project (Easter eggs), a once-present-but-now-past project (Big Screw), a once-future-once-present-but-now-past project (Cub Scout Decathlon), and a future project (Afterfinals, which will be at Camp Resolute).

The MVP reported her hosedness.

The AVP reported that she was spending too much time running around trying to arrange stuff for the PMD party we’re sponsoring.

The treasurer had nothing to report.

The fellowship chair asked everyone to attend Steak Fry, to PC I-Banquet, and to go to The Gondoliers and/or Little Shop of Horrors.

The APOcEd reported that the APOc existed and distributed them.

The historian reported that he did the SFF page for the scrapbook.

The PubDir had nothing to report.

The ASC reported (from afar) that she was “at a test (ugh) review. but if [she] was there, [she]’d actually have things to say, like: i cleaned up the office last week, and then *poof* a couple of days later it was darn hard to tell. it shouldn’t work this way. PLEEEEEEEEEEEZE clean up after yourselves… ”

The Chairman of the Advisory Committee had nothing to report.

The Big Screw PC reported that he hoped to get 2.70 to run.

The President reported that this was the first chapter meeting in three weeks, that hopefully replies for sectional service project would be showing up soon, that we have a fridge and a stereo system in the office, that as a one time exception ExecComm last week agreed to sponsor a PMD (People Making a Difference) fundraising party in Walker, and that he was willing to co-PC Afterfinals with a pledge.

As for new business, the following legislation was proposed for submission to Nationals:

Moved: That Item 7 under “Procedures” (The Ritual and Other Ceremonies) be amended to read:

Use of blindfolds is optional. Where chapters choose not to use blindfolds, or where state or local law or institutional regulations prohibit blindfolding, substitute closed eyes where appropriate.

Rationale: We feel that the use of physical blindfolds does not necessarily contribute anything to the effect of the ceremony which cannot be achieved equally well by simply having pledges close their eyes. Further, use of blindfolds is awkward and can be considered intimidating, while our asking the pledges (or neophytes) to close their eyes shows that we trust them.

The motion to submit this amendment to Nationals was carried many-0-1 by voice vote.

As for DSKs, the following people were nominated and discussed: Ezra Erb, Dave Cho, Kate Curtis, Bridget Spitznagel, Barbara Manganis, Crystal Reul, and Mike Person.

PFTGOTOs were as follows:

Amy: Thanks for making this issue of the APOc the biggest yet! Yale: Congrats to Kate on getting the fridge! Dave: Don’t send non-APO-specific email to apo-news. Henning: Thanks to 6.003, 1.00 and 5.60 for Big Screw.

ExecComm Summary: April 13, 1994

Start time was 7:32 pm. People were there. Reports that were not time-sensitive are as follows:

Historian (Dima): Still working on matching names to pictures.

Advisor (Joe): I have the certificates for everyone who took the CPR class.

Silkscreen (Dima): Rachel wants to put together a schedule of what runs people are planning to do.

President (Dave): I will be PCing I-Banquet. It will be at 100 Memorial Drive. Initiation will be on the third floor of building 26.

New business: For a full report on the facts of this business, please see Kevin or Dave who have written up a report.

Sometime between Thursday night and Friday morning, $97 disappeared from the strongbox, along with $6 from the treasurer’s drawer. There are lots of ways this could have happened; the key to the strong box was in the treasurer’s drawer, which is not secure.

Recommendations: Make sure the office is always secure and that the door is locked when there is no one conscious in the front part of the office. The missing money will be replaced from the executive discretionary fund so that Planned Parenthood (Big Screw) does not lose out. All receipts and money should now be deposited into the strongbox. No money should go into the treasurer’s drawer. The keys for the strongbox will be kept in the safe. A report was filed with the CPs. Possibilities: Talk to CAC about getting a retractor on the door. Repin the door lock so that it is more secure, but depends on how MIT feels about this.

PFTGOTOs:

Barbara: Richard and I have EMT exams on the 30th. Hannah: Premiere Orchestra concert, 2 pm, Saturday the 23rd. Dave: Drop Date Concert Choir concert, Verdi’s “Requiem,” Friday the 22nd, 8 pm in Kresge. Barbara: 6.111 still sucks. Dave: 6.003 sucks.

More Spring Breaking

Wait! Wait! Amy needs another 3 pages of material!

So I’ll write what I did during Spring Break.

Easy.

Friday, I did something, I’m sure. I think I slept.

Saturday, I think I had a netrek game.

Sunday, I think a bunch of us practiced ballroom dancing. Ballroom dancing is way cool. There are a bunch of different forms with totally different mindsets. Swing is a really casual and informal form. Rumba is a very sultry form; it’s not meant for seducing people, but just for being very expressive and very passionate. Waltz, of course, is very refined and very sophisticated. In its own way, it can be seductive, but it’s the kind of head-over-heels type of feeling. Tango requires a lot of self-confidence. It’s very different from Rumba, but I can’t really put my finger on what. That leaves Samba, which is another Latin dance, and is kinda a cross between Tango and Rumba: more passion than Tango, but faster than Rumba.

Monday I tooled 6.003. Ick!

Tuesday I tooled 6.003. Ick!

Wednesday I punted by stealing as much of ExecComm as possible for dinner. Afterwards, we went and saw Schindler’s List. That is such an amazing movie… everyone should definitely see it, but do not see it alone, and don’t forget to bring Kleenex. And don’t plan on doing anything afterwards.

Thursday was the Aquarium trip. That was fun. :)

Friday, Saturday, and Sunday I tooled. And slept.

That’s all. :)

dave

More Dave

Well, that last article certainly couldn’t have pushed everything to three pages, so here’s a few random annoucements:

MIT’s Concert Choir, Chamber Chorus, and Symphony Orchestra are performing Verdi’s “Requiem” this Friday at 8 pm in Kresge Auditorium. Tickets are $2 in advance for MIT students; $10 general admission. (Sorry; the music office is being fascist.) As of Tuesday’s rehearsal, we sound fantastic. Come listen to me, Richard Barbalace, Harry Tsai, and a couple hundred other people perform. It’s really cool. Besides, when else can you see Harry Tsai in a tuxedo?

Central Connecticut State University’s (CCSU) Epsilon Delta chapter is having their semi-normal [yes, that’s right, semi-normal] on Friday, May 13th. So, if you’re like me and are going to be done with finals by then and are interested in going, talk to me. No guarantees.

AfterFinals is still desperately seeking a PC. I’m already weighted down with I-Banquet. The rest of my ExecComm is drowning in work. I—we— really need a brother to step forward and deal with this. It’s not that hard…

Don’t forget, current officer hours are:

• Dave Cho: Sundays, noon-1 pm

• Kevin Dahm: Thursday nights. I don’t remember when, but if you look at the office door, you can see when it is.

Come on by if you have questions. We like visitors, really we do. :)

Hey, Amy, how much more do I need to write? :)

And One More Dave

OK, Amy needs one more page. Not a problem. I can spew on for hours. This is called wasting toner. It’s also called an exercise in stream of consciousness. Or, in this case, stream of unconsciousness. Not that I’m really that tired; I just feel that tired. But it’s still kinda fun, at least for me, to blather on and on and on.

Another announcement! They Might Be Giants will be performing at Harvard University’s Alumni Hall on Thursday, April 28, 1994 at 9:00 pm. Tickets are $16.00, and are available from both the Holyoke Information Center and the Sanders Theatre Box Office.

I guess it’s not too early to talk about the summer now, especially since I’m most certainly looking forward to it. For starters, Nancy Clark and I (and possibly John Grossi) will be working on a Summer Sectional Fellowship Event. Or at least, that was the idea at last time; it might be a service project. So we’ll be needing help with it (surprise!), so if you’re around during the summer, keep your eyes open for this.

New topic: On July 4th, it’s traditional that a lot of people descend on the Esplanade to listen to the Boston Pops perform and watch pretty fireworks. Assassins, SIPBions, and APOers convene in a traditional spot; watch your email for announcements at a later date.

The Minuteman Bike Trail connects between Alewife T Station and Belmont. However, there is another bike trail connecting Alewife and Porter. Therefore, if you bike down Hampshire street, you can, in relatively high safety, bike the fifteen miles between MIT and Belmont. On the way, you travel through Arlington and Lexington. People are already getting their bikes ready and starting to ramp up for summer. If you’re interested in joining this amoebic mob, send me mail.

Diskless is a group of people who play Ultimate Frisbee very poorly. I mean, we suck. Correction: we really suck. As member Chad Brown describes it, “We often know where we want the frisbee to go.” Currently, some subset of these people meets in Killian Court at 5:30 pm on Fridays to play. Come on by; if you suck, you’ll fit in perfectly. Or you can cheer; a lot of people just spectate.

OK, I’m afraid this’ll be too long, so let me wrap it up.

dave

Advertisements

WANTED: APOcrypha submissions of all shapes and sizes for the final issue of the term, premiering at I-Banquet, May 6. Must be two-dimensional and photocopiable. Titles will be awarded to distinguished contributors! Send your submission through email to

aihsu@mit.edu or apocrypha@mit.edu (whichever you feel like typ-

ing) or leave it in the APOcEd’s box in the APO office, or address it to Amy and pin it to the bulletin board, or address it to Amy and pin it on the office door, or address it to SPOO and leave it in the Technique office, or send it interdepartmental mail to Amy Hsu, New House, or send it U. S. mail to 476 Memorial Drive, Cambridge, MA 02139, or wave it in my face and say, “It’s for the APOc! Lose it and die!” or… well, you get the picture. Submit!

STILL HUH? Poor APOcEd is still having problems with page numbers. I LaTEXed the silly thing up to ten times and it didn’t seem to work. HELP!

The APOcrypha

Volume 80, Issue 6

Spring 1994ExecComm

President  Dave Cho  x5-6603  dcctdw
SVP  Bridget Spitznagel x5-6152  baspitz
MVP  Barbara Manganis x5-7407  autumn
AVP  Kate Curtis  628-5384 kcurtis
Treasurer  Kevin Dahm  628-5384 kdahm
Fellowship  Ann Yen  x5-7672  aiyen
APOcEd  Amy Hsu  x5-7678  aihsu
ASC  Traci Williams  x5-8455  tlw
PubDir  LeaAnn Collins  x5-8526  troth
Historian  Dima Dribinsky  x5-7413  dimon

Chair, Advisory Comm. Joe Kuchta

Brotherhood Expectations

Attend at least 6 service projects, including 2 on-campus and 2 off-campus Attend at least 2 fellowship events, including one of the major ones Attend at least 5 Chapter Meetings and 1 ExecComm Learn, teach, or practice a chapter skill Hold a leadership position within the chapter (e.g. PC, sub-PC, hold office) Attend the Pledge Term Review Meeting Attend ceremonies

Pledge Requirements

Attend at least 6 service projects, including 2 on-campus and 2 off-campus Attend at least 2 fellowship events, including one of the major ones Attend at least 2 Chapter Meetings and 1 ExecComm Attend at least two-thirds of scheduled Pledge Meetings Learn or teach a chapter skill Get a Big Brother within two weeks of pledging PC or sub-PC a service project or fellowship event Attend the Pledge Feedback Meeting